1st sum of money at 6%
2nd sum of money is twice the first sum + $1000 at 10% and the total annual income is $620..
how much is invested at each rate and how did you get it.. this has to be done in the i=prt form

Answer by scott8148(6628)

s=1st sum, so 2s+1000=2nd sum

(6%)(s)+(10%)(2s+1000)=620 __ .06s+.2s+100=620 __ .26s=520 __ s=2000

2s+1000=5000